OpenStudy (anonymous):

The area of a right triangle with one leg measuring 8 feet and hypotenuse measuring 17 feet is _____ square feet. 60 68 75

OpenStudy (anonymous):

Just use pythagurus.

OpenStudy (anonymous):

\[a^2 + b^2 =c^2\]

OpenStudy (anonymous):

where c= 17 b= 8

OpenStudy (anonymous):

solve for a

OpenStudy (anonymous):

\[289-64 = 225 \implies 15 \implies area = 4 \cdot 15 = 60\]

OpenStudy (zzr0ck3r):

yes 60 8^2 + x^2 = 17^2 so x = 15 then 1/2*8*15 we do not know or care which one is the base as long as we use the two sides that are not the hypotenuse for 1/2*b*h